OpenStudy (anonymous):

What is the distance between points (1, 4) and (9, 16)? Round to the nearest tenth. i got 14

OpenStudy (triciaal):

you can use pythagorean theorem the square of the distance = (change in y)^2 + (change in x)^2

rvc (rvc):

sqrtx2-x1+y2-y1 \[\sqrt{(9-1)^{2}+(16-4)^{2}}\] \[\sqrt{64+144} = \sqrt{208}\]

OpenStudy (triciaal):

to the nearest tenth means the answer to one decimal place

OpenStudy (anonymous):

so is 14.4??

OpenStudy (triciaal):

yes

OpenStudy (anonymous):

The midpoint between points (–3, 8) and (3, –4) is (0, y). What is y? can you help me

OpenStudy (triciaal):

always calculate to one more place and then round example for this problem you got 14.42 to one place 2 is less than 5 so drop it and the answer is 14.4

OpenStudy (triciaal):

if the calculated amount was 11.47 then the answer would be 11.5

OpenStudy (triciaal):

mid-point (x1 + x2)/2 , (y1 + y2)/2 use the absolute value

OpenStudy (triciaal):

how about if you draw it so it's easier to see?

OpenStudy (anonymous):



OpenStudy (anonymous):